OpenStudy (anonymous):

A right circular cone is inscribed in a cube having a diagonal which measures 10√3 cm. Find the volume of the cone and the volume of space between the cone and the cube.

OpenStudy (pawanyadav):

@ilikestudying140898 have a look what you wrote

OpenStudy (ilikestudying140898):

opps

OpenStudy (ilikestudying140898):

my bad..

OpenStudy (pawanyadav):

Diagonal of cube =√3 (side) And. From question. √3(side)=10√3 side=10 Now, diameter of cone=side of cube So radius =side/2=10/2=5

OpenStudy (pawanyadav):

Volume of cube =side^3=10^3=1000 cm^3 And volume of cone=1/3(πr^2h) Where r=5,h=10 So volume=?

OpenStudy (anonymous):

V=262

OpenStudy (pawanyadav):

Right

OpenStudy (anonymous):

what's the next?

OpenStudy (pawanyadav):

Volume of cube=?

OpenStudy (pawanyadav):

So volume of space between cube and cone= volume of cube- volume of cone

OpenStudy (pawanyadav):

=1000-262 cm^3 is the answer of second part.

OpenStudy (anonymous):

ahh okay. :) but I don't get this part "/2=10/2=5" from your second comment. hehe

OpenStudy (anonymous):

i get it now. thank you! :) this helps me a lot.